# Automated Federated Pipeline for Parameter-Efficient Fine-Tuning of Large Language Models

### v1

@@ Abstract.
Recently, there has been a surge in the development of advanced intelligent generative content (AIGC), especially large language models (LLMs). However, for many downstream tasks, it is necessary to fine-tune LLMs using private data. While federated learning offers a promising privacy-preserving solution to LLM fine-tuning, the substantial size of an LLM, combined with high computational and communication demands, makes it hard to apply to downstream tasks. More importantly, private edge servers often possess varying computing and network resources in real-world scenarios, introducing additional complexities to LLM fine-tuning. To tackle these problems, we design and implement an automated federated pipeline, named FedPipe, to fine-tune LLMs with minimal training cost but without adding any inference latency. FedPipe firstly identifies the weights to be fine-tuned based on their contributions to the LLM training. It then configures a low-rank adapter for each selected weight to train local low-rank adapters on an edge server, and aggregate local adapters of all edge servers to fine-tune the whole LLM. Finally, it appropriately quantizes the parameters of LLM to reduce memory space according to the requirements of edge servers. Extensive experiments demonstrate that FedPipe expedites the model training and achieves higher accuracy than state-of-the-art benchmarks.
@@
federated learning, large-scale language model, fine-tune.
The major part of this work was done when Zihan Fang worked as master student at Fudan University, with the remaining part completed during her PhD studies at City University of Hong Kong.
ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Machine learning
1. Introduction
Figure 1. A scenario of LLM FL via PEFT with edge servers.
@@ 1. Introduction
We are now witnessing the tremendous success of advanced intelligent generative content (AIGC), especially large language models (LLMs), such as OpenGPT (Radford et al., 2019; Brown et al., 2020), LLaMa (Touvron et al., 2023), and Palm (Chowdhery et al., 2023), in our daily life. Due to the huge number of model parameters (e.g., OpenGPT3 has 175 B parameters), those LLMs deliver outstanding performance on various natural language processing (NLP) applications, especially, questions and answers (Jiang et al., 2023), sentiment analysis (Englhardt et al., 2023; Nan et al., 2021), etc. It is believed that many new exciting applications will come out in the near future such as smart healthcare, intelligent transportation, and image classification (Huang et al., 2023; Hu et al., 2023a; Yuan et al., 2023; Hu et al., 2024). Considering the state-of-the-art workflow of LLMs, generally, we can summarize into three phases: 1) pre-training LLMs from scratch with extensive computing resources (e.g., training GPT3-1.3B model requires 64 Tesla V100 GPUs for one week (Yuan et al., 2022)) and tremendous text corpora (e.g., Wikipedia); 2) fine-tuning pre-trained LLMs for various downstream tasks; and 3) deploying fine-tuned LLMs at an edge server to infer from input data.
@@
Due to privacy concerns, even if many parties with the same downstream task demand a shared LLM, they cannot contribute their own data to fine-tune it (Lin et al., 2023b). For example, several hospitals would like to fine-tune an LLM for healthcare via their edge servers (Karargyris et al., 2023), but they cannot share patients’ data. Although federated learning (FL) is a promising solution to address this privacy issue (Shin et al., 2022; Panchal et al., 2023; Dai et al., 2023a; Shuai et al., 2022; Dai et al., 2023b), it is non-trivial to fine-tune LLMs via FL due to the limited computing capability of edge servers. In typical configurations, edge servers are outfitted with low-cost, commercially available GPUs, such as the NVIDIA GeForce RTX 3090-Ti. These GPUs, while powerful for many applications, are outmatched by the more robust GPUs found in data center environments (Panopoulos et al., 2023; Cai et al., 2023a). Specifically, data centers often deploy data-center-grade GPUs like the NVIDIA A100 and V100, characterized by significantly higher communication bandwidth (e.g., NVLink), larger graphic memory, and enhanced computing capabilities. This distinction in hardware resources has critical implications for computational tasks (Chen et al., 2023; Ouyang et al., 2023b, a), notably, edge servers face challenges in supporting full parameter fine-tuning of LLMs within standard FL frameworks.
@@
Fortunately, the integration of FL and Parameter-Efficient Fine-Tuning (PEFT) may solve the above challenges. Recent research efforts (Houlsby et al., 2019; Pfeiffer et al., 2021; Hu et al., 2022; Karimi Mahabadi et al., 2021; Sung et al., 2021; Zhang et al., 2023a; Hu et al., 2023b; Sheng et al., 2023) have explored PEFT methods, in particular, ADAPTER (Houlsby et al., 2019; Pfeiffer et al., 2021; Karimi Mahabadi et al., 2021) and LoRA (Hu et al., 2022; Sheng et al., 2023), to mitigate the challenges of intensive computational resources required by LLM fine-tuning. Figure 1 illustrates how to integrate PEFT with FL for LLMs. The central server distributes a pre-trained LLM to edge servers (participants), and then edge servers perform local training using PEFT methods. After training, the central server aggregates these updates and redistributes the enhanced LLM to edge servers for subsequent training rounds. This iteration continues until the LLM achieves a predefined accuracy threshold. Recent pioneer work, FedAdapter (Cai et al., 2023b), adopts ADAPTER to achieve efficient FL for NLP downstream fine-tuning, but mainly focuses on standard language model, such as BERT (Kenton and Toutanova, 2019). A few research works (Liu et al., 2023; Che et al., 2023) study LLM FL with PEFT, but they always assume every edge server with homogeneous resources (i.e., computing power and memory). Consequently, these frameworks are not well adapted to heterogeneous environments and cause deterioration in training performance. In practice, it is still non-trivial to integrate PEFT with FL in LLMs for downstream tasks.
@@
In this paper, we propose and implement an automated federated pipeline (FedPipe) for LLM FL fine-tuning. Unlike existing PEFT methods, FedPipe does not aim to introduce new PEFT techniques. Instead, it serves as a foundational framework to enhance any existing PEFT method. This is achieved by systematically decoupling, abstracting, and automating the operations and parameters of these methods. As a result, application and algorithm developers can efficiently outsource the implementation intricacies of LLM FL fine-tuning, allowing them to concentrate on their main services or model design. Despite the promising potential of FedPipe, how to implement this framework poses significant challenges. First, given the varying computing resources across edge servers, this heterogeneity causes a pronounced straggler problem in FL (Reisizadeh et al., 2022; Wu et al., 2021; Zhang et al., 2023b; Lin et al., 2023a; Horvath et al., 2021; Lyu et al., 2023; Ouyang et al., 2022). For instance, two edge servers with slightly different GPU memories (e.g., 22 GB vs. 24 GB) can experience a training time discrepancy of several hours per communication round in FL. Second, to offer better fine-tuning performance, FedPipe enables each edge server to identify and prioritize important weights for adapter construction. However, assessing and modeling the importance of these weights remains an open problem. Finally, each adapter within FedPipe must align well with the computing budget of its respective edge server, which demands careful adjustment of model parameters. These challenges underscore the complexity and necessity of careful design of FedPipe.
@@
In this paper, we first conduct pilot measurements to analyze the above challenges and acquire the corresponding insights to guide the system design of FedPipe. Then, we formulate the automated LLM FL fine-tuning pipeline as a Mixed-Integer Linear Programming (MILP) optimization problem. To solve the MILP problem, we propose a two-stage method, comprised of trainable weights identification method, and fast search algorithm. The adaptive identification method of important weights is used to figure out the best adapter of an LLM to fine-tune based on their importance score for each edge server. The method parameterizes incremental updates (i.e., trainable weights) using singular value decomposition (SVD) to effectively prune unimportant updates with the lower singular values, leading to constructing the efficient adapter with less trainable weights. According to the adapters of edge servers, by considering heterogeneous computing resources of edge servers, we design a fast search algorithm to select the best set of rank and batch size for local training at each edge server. Finally, to consider diverse memory budgets for edge servers, and further, accelerate their training speed, we leverage a simple quantization-agnostic backward pass to adaptively achieve low-precision LLM weights from a custom black-box quantization module. Moreover, different from existing LLM FL aggregation approaches, we design a lightweight partial weights aggregation approach by only aggregating local adapters to improve communication efficiency.
We summarize the main contributions of FedPipe in the following:
@@
  • To the best of our knowledge, FedPipe is the first automated federated pipeline of fine-tuning LLMs for all kinds of downstream tasks.
@@
  • We model the pipeline as an MILP optimization problem and design an efficient algorithm to solve it.
@@
  • By taking heterogeneous computing resources at edge servers into consideration, we develop an effective method to find the best low-rank adaptation structure for each edge server, which is needed in efficiently solving our MILP problem.
@@
  • To further reduce the computation in solving our MILP at edge servers, we also propose a fast search algorithm to dynamically identify important parameters for local adapter training.
@@
  • Furthermore, according to those parameters, and memory budgets of edge servers, we quantize the local models into different quantization bits to increase training efficiency.
@@
Figure 2. An illustration of LoRA method.
The rest of this paper is organized as follows. Section 2 introduces the background and motivation and the system design of FedPipe is presented in Section 3. Sections 4 and 5 respectively elaborate FedPipe’s implementation and report the extensive evaluations of FedPipe. Section 6 briefly discusses the related works, followed by the conclusion in Section 7.
2. Background and Motivation
@@
Figure 3. The training time, and exchanged parameters of LoRA and full parameter fine-tuning for GPT-2, where $r$ is the rank size of LoRA adapter.
2.1. Injecting PEFT into Federated Learning on LLMs
Due to the constrained computing resources at an edge server (i.e., commercial GPUs, such as NVIDIA GeForce RTX 3090-Ti) and the large number of parameters in an LLM, full parameter fine-tuning (i.e., initializing the model with pre-trained weights, updating all parameters, and generating separate instances for various tasks) at each edge server leads to unacceptable computing and networking latency in both model training and aggregation. Table 1 illustrates the number of parameters and sizes of several popular LLMs. Obviously, the smallest model, GPT-2 (Radford et al., 2019) with 774 M parameters, is over 7 times larger than the conventional language model (e.g., BERT (Kenton and Toutanova, 2019)) with 110 M parameters.
@@ 2.1. Injecting PEFT into Federated Learning on LLMs
Table 1. Popular LLMs
@@
To combat the above stumbling obstacles in LLM FL, PEFT employed in downstream task training is a potential solution. Different from the full parameter fine-tuning, PEFT only fine-tunes parts of neurons to reduce trainable parameters in a model, leading to less computing cost. Recently, there are mainly two types of PEFT, ADAPTER (Houlsby et al., 2019; Pfeiffer et al., 2021; Karimi Mahabadi et al., 2021) and LoRA (Hu et al., 2022; Sheng et al., 2023). As shown in Figure 2, ADAPTER inserts a few layers for each transformer block (Houlsby et al., 2019) and only trains those layers in a pre-trained model. Apparently, ADAPTER introduces extra latency in the inference phase, due to its additional layers (Hu et al., 2022). However, LoRA is an adapter with low-rank adaptation without additional inference latency that injects trainable rank decomposition matrices into each transformer layer, but freezes all weights of a pre-trained model. Many works (Hu et al., 2022, 2023b; Sheng et al., 2023) have verified that LoRA outperforms ADAPTER, and hence in the following, we only focus on LoRA.
@@
To better understand the computing and communication bottlenecks in the LLM FL, we set up two experiments using one of the most prevalent LLMs, GPT-2. Figure LABEL:fig:llm_train:computing and LABEL:fig:llm_train:communication show the training time for a batch of data, and exchanged parameters of full parameter fine-tuning and PEFT. Full parameter fine-tuning obtains more than 1.4 and 1000 times larger than those of PEFT in training and exchanging, respectively. Apparently, instead of the entire model fine-tuning, PEFT can significantly mitigate computing and communication overhead. However, we also meet three main challenges to integrate PEFT into LLM FL. We will use “training” to represent fine-tuning in the following sections.
Figure 4. The straggler problem on GPT-2 and BERT.
2.2. Severe Straggler Problem in LLM FL
@@
State-of-the-art LLM FL frameworks (Liu et al., 2023; Che et al., 2023) usually assume that each edge server has sufficient local resources (e.g., computing power, communication bandwidth, and memory space) to perform local model training. Therefore, a central server can easily aggregate these local models to achieve good performance. However, in practice, the available resources of various edge servers vary significantly, and the allocated resources for training may change at run-time depending on how on-demand running programs prioritize resource allocation. More importantly, due to the large parameters of LLM, the resource heterogeneity of edge servers leads to significantly different training time, and a severe straggler problem in model aggregation.
@@
To better understand the impact of heterogeneous resources on the straggler problem for LLM FL, we conduct two experiments for a standard model and LLM. We integrate LoRA into FedAvg (McMahan et al., 2017), one of the most prevalent FL frameworks, to evaluate its performance on the widely recognized LLM GPT-2 and standard model BERT for homogeneous and heterogeneous cases. We set up two PC servers (i.e., edge servers) with/without varying computing power by controlling different graphic memories of GPU (22 GB and 24 GB), but the same ones (24 GB) for homogeneous and heterogeneous cases, respectively. The result is shown in Figure 4 where the under-training rate is defined by the number of under-training edge servers divided by the total number of edge servers and average aggregation time means the per-round time from the start of local model training for a batch of data to the end of model aggregation.
@@
Figure 5. GPT-2 training performance with different trainable weights, where the trainable weights combinations corresponding to quantities 1, 2, 3, and 4 are $\left\{{{{\bf{W}}_{q}}}\right\}$, $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}}}\right\}$, $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}},{{\bf{W}}_{v}}}\right\}$, and $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}},{{\bf{W}}_{v}},{{\bf{W}}_{o}}}\right\}$, respectively.
Figure 6. The PPL at the $50$-th and $70$-th training rounds and the STDs of the PPL across 100 training rounds.
2.3. Weight-Level Configuration

### v2

@@ Abstract
Recently, there has been a surge in the development of advanced intelligent generative content (AIGC), especially large language models (LLMs). For many downstream tasks, it is necessary to fine-tune LLMs using private data. While federated learning offers a promising privacy-preserving solution to LLM fine-tuning, the substantial size of an LLM, combined with high computational and communication demands, makes it hard to apply to handle downstream tasks. More importantly, private edge servers often possess varying computing and network resources in real-world scenarios, introducing additional complexities to LLM fine-tuning. To tackle these problems, we design and implement an automated federated pipeline, named FedPipe, to fine-tune LLMs on heterogeneous edge servers with minimal training cost and no additional inference latency. FedPipe firstly identifies the weights to be fine-tuned based on their contributions to the LLM training. It then configures a low-rank adapter for each selected weight within the resource constraints of the edge server and aggregates these local adapters from all edge servers to fine-tune the whole LLM. Finally, it appropriately quantizes the parameters of LLM to reduce memory space according to the requirements of edge servers. Extensive experiments demonstrate that FedPipe expedites model training and achieves higher accuracy than the state-of-the-art benchmarks.
Index Terms:
@@
federated learning, large-scale language model, fine-tuning.
I Introduction
@@
We are now witnessing the tremendous success of advanced intelligent generative content (AIGC) [1], especially large language models (LLMs), such as OpenGPT [2, 3], LLaMa [4], and Palm [5], in our daily life. Due to the huge number of model parameters (e.g., OpenGPT3 has 175 B parameters), those LLMs deliver outstanding performance on various natural language processing (NLP) applications, especially, questions and answers [6, 7, 8], sentiment analysis [9, 10], etc. It is believed that many new exciting applications will come out in the near future such as smart healthcare, intelligent transportation, and image classification [11, 12, 13, 14, 15, 16]. Considering the state-of-the-art workflow of LLMs, generally, we can summarize into three phases: 1) pre-training LLMs from scratch with extensive computing resources (e.g., training GPT3-1.3B model requires 64 Tesla V100 GPUs for one week [17]) and tremendous text corpora (e.g., Wikipedia); 2) fine-tuning pre-trained LLMs for various downstream tasks; and 3) deploying fine-tuned LLMs at an edge server to infer from input data.
@@
Due to privacy concerns, even if many parties with the same downstream task demand a shared LLM, they cannot directly contribute their raw data to fine-tune it [18, 19, 20, 21]. For example, several hospitals would like to fine-tune an LLM for a special disease diagnosis via their edge servers [22, 23], but they cannot share their patients’ data. Although federated learning (FL) is a promising solution to address this privacy issue [24, 25, 26, 27, 28, 29], it is non-trivial to fine-tune LLMs via FL due to the limited computing capability of edge servers. In typical configurations, edge servers are outfitted with low-cost, commercially available GPUs, such as the NVIDIA GeForce RTX 3090-Ti. These GPUs, while powerful for many applications, are outmatched by the more robust GPUs found in data center environments [30, 31]. Specifically, data centers often deploy data-center-grade GPUs like the NVIDIA A100 and V100, characterized by significantly higher communication bandwidth (e.g., NVLink), larger graphic memory, and enhanced computing capabilities. This distinction in hardware resources has critical implications for computational tasks [32, 33, 34], notably, edge servers face challenges in supporting full parameter fine-tuning of LLMs within standard FL frameworks.
Figure 1: A scenario of LLM FL via PEFT with edge servers.
@@
Fortunately, the integration of FL and Parameter-Efficient Fine-Tuning (PEFT) may solve the above challenges. Recent research efforts [35, 36, 37, 38, 39, 40, 41, 42] have explored PEFT methods, in particular, ADAPTER [35, 36, 38] and LoRA [37, 42], to mitigate the challenges of intensive computational resources required by LLM fine-tuning. Fig. 1 illustrates how to integrate PEFT with FL for LLMs. The central server distributes a pre-trained LLM to edge servers (participants), and then edge servers perform local training using PEFT methods. After training, the central server aggregates these updates and redistributes the enhanced LLM to edge servers for subsequent training rounds. This iteration continues until the LLM achieves a predefined accuracy threshold. Recent work, FedAdapter [43], adopts ADAPTER to achieve efficient FL for NLP downstream fine-tuning, but mainly focuses on standard language model, such as BERT [44]. A few research works [45, 46] study LLM FL with PEFT, but they always assume every edge server with homogeneous resources (i.e., computing power and memory). Consequently, these frameworks are not well adapted to heterogeneous environments and cause deterioration in training performance. In practice, it is still non-trivial to integrate PEFT with FL in LLMs for downstream tasks.
@@
In this paper, we propose and implement an automated federated pipeline (FedPipe) for LLM FL fine-tuning. Unlike existing PEFT methods, FedPipe does not aim to introduce new PEFT techniques. Instead, it serves as a foundational framework to enhance any existing PEFT method. This is achieved by systematically decoupling, abstracting, and automating the operations and parameters of these methods. As a result, application and algorithm developers can efficiently outsource the implementation intricacies of LLM FL fine-tuning, allowing them to concentrate on their main services or model design. Despite the promising potential of FedPipe, how to implement this framework poses significant challenges. First, given the varying computing resources across edge servers, this heterogeneity causes a pronounced straggler problem in FL [47, 48, 49, 50, 51, 52]. For instance, two edge servers with slightly different GPU memories (e.g., 22 GB vs. 24 GB) can experience a training time discrepancy of several hours per communication round in FL. Second, to offer better fine-tuning performance, FedPipe enables each edge server to identify and prioritize important weights for adapter construction. However, assessing and modeling the importance of these weights remains an open problem. Finally, each adapter within FedPipe must align well with the computing budget of its respective edge server, which demands careful adjustment of model parameters. These challenges underscore the complexity and necessity of careful design of FedPipe.
@@
In this paper, we first conduct pilot measurements to analyze the above challenges and acquire the corresponding insights to guide the system design of FedPipe. Then, we formulate the automated LLM FL fine-tuning pipeline as a Mixed-Integer Linear Programming (MILP) optimization problem. To solve the MILP problem, we propose a two-stage method, comprised of trainable weights identification method and fast search algorithm. The adaptive identification method of important weights is used to figure out the best adapter of an LLM to fine-tune based on their importance score for each edge server. This method parameterizes incremental updates (i.e., trainable weights) using singular value decomposition (SVD) to effectively prune unimportant updates with the lower singular values, leading to the efficient adapter with fewer trainable weights. According to the adapters of edge servers, by considering heterogeneous computing resources of edge servers, we design a fast search algorithm to select the best set of rank and batch size for local training at each edge server. Finally, to consider diverse memory budgets for edge servers, and further accelerate their training speed, we leverage a simple quantization-agnostic backward pass to adaptively achieve low-precision LLM weights from a customized black-box quantization module. Moreover, different from existing LLM FL aggregation approaches, we design a lightweight partial weights aggregation approach by only aggregating local adapters to improve communication efficiency.
We summarize our main contributions of FedPipe as follows.
@@
  • FedPipe is the first general automated federated pipeline of fine-tuning LLMs for all kinds of downstream tasks.
@@
  • We formulate the pipeline as an MILP optimization problem and develop an efficient algorithm to solve it.
@@
  • By taking heterogeneous computing resources at edge servers into consideration, we develop an effective method to find different low-rank adapter structures for heterogeneous edge servers.
@@
  • To further reduce the computational complexity in solving our MILP at edge servers, we also design a fast search algorithm to dynamically identify important parameters for local adapter training.
@@
  • Furthermore, according to those parameters and memory budgets of edge servers, we quantize the local models into different quantization bits to increase training efficiency.
@@
The rest of this paper is organized as follows. Section II introduces the background and motivation and the system design of FedPipe is presented in Section III. Sections IV and V respectively elaborate FedPipe’s implementation and report the extensive evaluations of FedPipe. Section VI briefly discusses the related works, followed by the conclusion in Section VII.
II Background and Motivation
@@
II-A Injecting PEFT into Federated Learning on LLMs
Figure 2: An illustration of LoRA method.
Due to the constrained computing resources at an edge server (i.e., commercial GPUs, such as NVIDIA GeForce RTX 3090-Ti) and the large number of parameters in an LLM, full parameter fine-tuning (i.e., initializing the model with pre-trained weights, updating all parameters, and generating separate instances for various tasks) at each edge server leads to unacceptable computing and networking latency in both model training and aggregation. Table I illustrates the number of parameters and sizes of several popular LLMs. Obviously, the smallest model, GPT-2 [2] with 774 M parameters, is over 7 times larger than the conventional language model (e.g., BERT [44]) with 110 M parameters.
@@ II-A Injecting PEFT into Federated Learning on LLMs
TABLE I: Popular LLMs
@@
To combat the above stumbling obstacles in LLM FL, PEFT employed in downstream task training is a potential solution. Different from the full parameter fine-tuning, PEFT only fine-tunes parts of neurons to reduce trainable parameters in a model, leading to less computing cost. Recently, there are mainly two types of PEFT, ADAPTER [35, 36, 38] and LoRA [37, 42]. As shown in Fig. 2, ADAPTER inserts a few layers for each transformer block [35] and only trains those layers in a pre-trained model. Apparently, ADAPTER introduces extra latency in the inference phase, due to its additional layers [37]. However, LoRA is an adapter with low-rank adaptation without additional inference latency that injects trainable rank decomposition matrices into each transformer layer, but freezes all weights of a pre-trained model. Many works [37, 41, 42] have verified that LoRA outperforms ADAPTER, and hence in the following, we only focus on LoRA.
@@
To better understand the computing and communication bottlenecks in the LLM FL, we set up two experiments using one of the most prevalent LLMs, GPT-2. Fig. 3a and 3b show the training time for a batch of data and exchanged parameters of full parameter fine-tuning and PEFT. Full parameter fine-tuning obtains more than 1.4 and 1000 times larger than those of PEFT in training and exchanging, respectively. Apparently, instead of the entire model fine-tuning, PEFT can significantly mitigate computing and communication overhead. However, we also meet three main challenges to integrate PEFT into LLM FL. We will use “training” to represent fine-tuning in the following sections.
(a) Computing overhead
(b) Communication overhead
Figure 3: The training time, and exchanged parameters of LoRA and full parameter fine-tuning for GPT-2, where $r$ is the rank size of LoRA adapter.
(a) Under-training rate
(b) Average aggregation time
Figure 4: The straggler problem on GPT-2 and BERT.
II-B Severe Straggler Problem in LLM FL
@@
State-of-the-art LLM FL frameworks [45, 46] usually assume that each edge server has sufficient local resources (e.g., computing power, communication bandwidth, and memory space) to perform local model training. Therefore, a central server can easily aggregate these local models to achieve good performance. However, in practice, the available resources of various edge servers vary significantly, and the allocated resources for training may change at run-time depending on how on-demand running programs prioritize resource allocation. More importantly, due to the large parameters of LLM, the resource heterogeneity of edge servers leads to significantly different training time, and a severe straggler problem in model aggregation.
@@
To better understand the impact of heterogeneous resources on the straggler problem for LLM FL, we conduct two experiments for a standard model and LLM. We integrate LoRA into FedAvg [53], one of the most prevalent FL frameworks, to evaluate its performance on the widely recognized LLM GPT-2 and standard model BERT for homogeneous and heterogeneous cases. We set up two PC servers (i.e., edge servers) with/without varying computing power by controlling different graphic memories of GPU (22 GB and 24 GB), but the same ones (24 GB) for homogeneous and heterogeneous cases, respectively. The result is shown in Fig. 4 where the under-training rate is defined to be the number of under-training edge servers divided by the total number of edge servers and the average aggregation time means the per-round time from the start of local model training for a batch of data to the end of model aggregation.
@@
Figure 5: GPT-2 training performance with different trainable weights, where the trainable weights combinations corresponding to quantities 1, 2, 3, and 4 are $\left\{{{{\bf{W}}_{q}}}\right\}$, $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}}}\right\}$, $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}},{{\bf{W}}_{v}}}\right\}$, and $\left\{{{{\bf{W}}_{q}},{{\bf{W}}_{k}},{{\bf{W}}_{v}},{{\bf{W}}_{o}}}\right\}$, respectively.
(a) PPL
(b) STD of PPL
Figure 6: The PPL at the $50$-th and $70$-th training rounds and the STDs of the PPL across 100 training rounds.
II-C Weight-Level Configuration